1. points are (-4,7) and (2,x) m=-1/3, find for X
2. 1/2x-3/4y=-4
1/4x+3/2y=13 solve using addition method
3. 2/3x-1/6y=0
4/3x+5/6y=14 solve using addition method
4. 1/2x+1/3y=-13/6
4/5x + 5/6y=9/10 solve using addition method

Answer by lynnlo(4176)

(1)x=1/282m
(2)y=8,x=4
(3)y=12,x=3
(4)y=131/9,x=-379/27